### v2

# Uniform approximation of Betti numbers

Abstract.
We prove that Lefschetz’s principle of approximating the cohomology of a possibly singular affine scheme of finite type over a field by the cohomology of a suitable (thickening of a) hyperplane section can be made uniform: in the affine case, we can choose the hyperplane section independently of the cohomology. Using Jouanolou’s trick, this gives a new way to bound the Betti numbers of quasi-projective schemes over a field, independently of the cohomology. This is achieved through a motivic version of Deligne’s generic base change formula and an axiomatic presentation of the theory of perverse sheaves. These methods produce generating families of Voevodsky motivic sheaves that are realized in perverse sheaves over any base of equal characteristic.
Introduction
Since Grothendieck introduced étale cohomology, proving that the Betti numbers of $\ell$-adic cohomology of algebraic varieties over a separably closed field are independent of $\ell$ remains a deep open problem in positive characteristic. The case of smooth and proper algeraic varieties is reached by weight techniques, using Deligne’s proof of the Weil conjectures, but there is not any other general result in this direction. The next best thing consists in providing uniform bounds of Betti numbers. This can be done by dévissage from the case of smooth and proper algebraic varieties, and then applying descent techniques together with de Jong’s alteration theorem, as explained by Katz [Kat01, Theorem 5] who credits the result to de Jong and Berthelot. There are results for non-constant coefficients as well, working with compatible families [KL85, Kat01, Org19].
The present notes give new proofs of existence of such uniform bound, using motivic techniques (in particular, we do not only prove independence of $\ell$ results but rather independence against a large class of Weil cohomologies). We apply Lefschetz’s principle of approximating a quasi-projective scheme $X$ by a suitable hyperplane section. Of course, this needs adjustments in order to take into account the singularities of $X$ as well as the fact that we want the choice of the hyperplane section to be independent of the cohomology we work with (e.g. $\ell$-adic cohomology of schemes of finite type over a separably closed field $k$, with $\ell$ distinct from the characteristic of $k$, or Berthelot’s rigid cohomology if $k$ is perfect of characteristic $p>0$).
To be more precise, the general strategy goes as follows. In this introduction, in order to simplify things, we work over an infinite field $k$ of characteristic $p$. Using Jouanolou’s trick [Jou73, Lemme 1.5], since the formation of Voevodsky motives is $\mathbf{A}^{1}$-homotopy invariant, the motive of any quasi-projective scheme is isomorphic to the motive of an affine one. Therefore, we can focus on affine schemes of finite type over $k$. For such a scheme $X$ of dimension $d$, one defines the notion of Leschetz subscheme: a closed subscheme $Z$ of $X$ of dimension d1d-1, with affine complement, that contains all irreducible components of $X$ of dimension <d<d and the singular locus of $X$, as well as a hyperplane section satisfying a suitable transversality condition, expressed in the language of the six operations for motivic sheaves; see Definition 3.2 below. We define the motive of the pair $(X,Z)$ through the cofiber sequence
$M(Z)\to M(X)\to M(X,Z)$
in Voevodsky’s category $\mathit{DM}_{c}(\mathrm{Spec}(k))$ (of constructible étale motives). We prove that, for any prime number $\ell\neq p$, the $\ell$-adic realization of $M(X,Z)$ is concentrated in degree d-d and that its unique non-trivial cohomology group is a free $\mathbf{Z}_{\ell}$-module of finite type, using the theory of $\ell$-adic perverse sheaves [^1] This is where the fact that we work with affine schemes matters: the main tool is then Artin’s vanishing theorem in order to prove Beilinson’s transversality Theorem 2.22 below.; see Lemma 3.4. Since the $\ell$-adic realization of motivic sheaves is symmetric monoidal, it is compatible with the formation and the computation of traces of endomorphisms of constructible motives, in particular of Euler characteristics (these are traces of identities). In this case, this means that the Euler characteristic of M(X,Z)[d]M(X,Z)[-d], seen as a motive, coincides with the rank $\mathit{rk}(X,Z)$ of the unique non-trivial cohomology group of its $\ell$-adic realization, proving that the latter is independent of $\ell$. This yields isomorphisms
$H^{q}_{\mathrm{\acute{e}t}}(\bar{X},\mathbf{Z}_{\ell})\cong H^{q}_{\mathrm{\acute{e}t}}(\bar{Z},\mathbf{Z}_{\ell})\quad\text{for $q\neq d,d-1$,}$
together with an exact sequence
$0\to H^{d-1}_{\mathrm{\acute{e}t}}(\bar{X},\mathbf{Z}_{\ell})\to H^{d-1}_{\mathrm{\acute{e}t}}(\bar{Z},\mathbf{Z}_{\ell})\to\mathbf{Z}_{\ell}^{\mathit{rk}(X,Z)}\to H^{d}_{\mathrm{\acute{e}t}}(\bar{X},\mathbf{Z}_{\ell})\to 0\,.$
In particular we get a bound of the $d$th Betti number of $X$ by the motivically defined non-negative integer $\mathit{rk}(X,Z)$, and we observe that the Betti numbers of $X$ in lower degree are bounded by those of $Z$. This is meaningful because, following ideas from Beilinson [Bei87], we apply a motivic variant of Deligne’s generic base change theorem, to deduce a motivic shadow of the weak Lefschetz theorem (Theorem 1.16 below). This implies that, for any closed subscheme $Y\subset X$ of dimension <d<d, there is a Lefchetz subscheme $Z\subset X$ containing $Y$ (this is where we need $k$ to be infinite: in fact, we prove that the family of eligible hyperplane sections contains the rational points of an open dense subscheme of the space of all hyperplane sections); see Prop. 3.3. Therefore, following a longstanding tradition in Algebraic Topology [Mil64, Tho65], we can produce a Lefschetz filtration of $X$: a finite sequence of closed subschemes of the form
$\varnothing=X_{-1}\subset X_{0}\subset\ldots\subset X_{d-1}\subset X_{d}=X$
where, for $0\leq i\leq d$, each $X_{i-1}$ is a Lefschetz subscheme of $X_{i}$. This leads to a filtration of the motive of $X$ whose associated graded object yields a complex with terms of the form
$M(X_{i},X_{i-1})[-i]\,.$
This means that the $i$th $\ell$-adic Betti number of the affine scheme $X$ (both $\mathbf{Q}_{\ell}$- or with $\mathbf{F}_{\ell}$-coefficients) is bounded by $\mathit{rk}(X_{i},X_{i-1})$ for all $\ell\neq p$; see Theorem 3.6.
However, our proof does more than providing a bound that is independent of $\ell$: it provides a uniform bound that is valid in any realization of Voevodsky’s motivic sheaves that has a theory of perverse sheaves (including the properties that Verdier duality induces an equivalence on constructible objects and that Artin’s vanishing theorem holds); see Definition 2.7. By developing this axiomatic approach, we also try to reach a wide enough level of generality, geometrically: instead of working over a field, we develop enough material to prove all the main ingredients over a scheme of finite type over a field: what is true over an infinite field is in fact true generically, possibly after pulling back along a flat finite cover of a dense open subset of any algebraic variety (from this point of view, we can work over any field $k$). Because what is at stake, after all, is the problem of finding big enough families of motivic sheaves that are known to be realized into ordinary sheaves or into perverse sheaves. This is what we do at the end of this article (Prop. 3.9 and 3.17), thus providing constructions of $t$-structures on $\mathit{DM}(X,\mathbf{Q})$ that are candidates for the perverse motivic $t$-structure (3.18). These constructions, by finding nice generators for categories of motivic sheaves, may be seen as a way to provide uniform bounds of Betti numbers of all sorts, not only for the cohomology of algebraic varieties, but for the cohomology of any constructible motivic sheaf (this includes cohomology with compact support or not, as well as nearby cycles, thanks to Ayoub’s work [Ayo07]), possibly in relative situations. Finding explicit generators for such $t$-structures is also part of the standard tools to develop the theory of Nori motives and to relate them with Voevodsky motives in characteristic zero. Therefore, the results of the present text open the way to a uniform theory of Nori motives [HMS17, Ara05, IM19], possibly over any field: instead of working with a single cohomology theory, we can now consider the family of all cohomology theories appearing in a six-functors formalism in which there is a good notion of perverse sheaves. This is strongly related with Ayoub’s constructions [Ayo23] of new Weil cohomologies in positive characteristic and their associated Hopf algebroids.
1. Motivic shadow of the weak Lefschetz Theorem
$1.$
We consider a $\mathbf{Z}$-linear six-functors formalism $D$ in the sense of We assume further more that, for any $X$, there is a full stable subcategory $D_{c}(X)$ of $D(X)$, whose objects are called constructible objects with the following properties:
  • (i)
    $D_{c}(X)$ contains the $\otimes$-unit $\mathbf{1}_{X}$, is closed under tensor product, under pullbacks of the form $f^{*}$ for any map $f$, and under direct image with compact support $f_{!}$ for any separated morphism of finite presentation $f$;
  • (ii)
    for any $X$ and any $F$ in $D_{c}(X)$, there is a dense open subscheme $U\subset X$ such that $F_{|U}$ is dualizable (in the sense of [Lur17, Section 4.6.1]);
  • (iii)
    for any $X$, any dualizable object of $D(X)$ is constructible;
  • (iv)
    constructible objects form an hypersheaf of $\infty$-categories for the étale topology.
Remark that, in the case of schemes with noetherian underlying space, these properties determine $D_{c}(X)$ uniquely.
Example 1.2.
Let $\mathit{DM}(X)=\mathrm{Ind}(\mathit{DM}_{h,lc}(X))$ the $\infty$-category of ind-objects on the $\infty$-category of locally constructible objects of the $\infty$-category of $h$-motives in the sense of [CD16, Def. 6.3.1]. Then $\mathit{DM}$ satisfies all the properties above, at least if we restrict to noetherian schemes (but this is not even necessary); this follows easily from see [CD16, Theorem 6.3.26]. For any commutative ring $\Lambda$, there is a $\Lambda$-linear version $\mathit{DM}(X,\Lambda)$ defined as the $\Lambda$-modules in $\mathit{DM}(X)$.
If $\Lambda$ is a commutative ring of characteristic $n>0$, there is a rigidity theorem in the following sense: for $n$ prime to the residue characteristics of $X$, there is a canonical equivalence between compact $\Lambda$-modules in $\mathit{DM}(X)$ and ordinary constructible sheaves of $\Lambda$-modules of finite tor-dimension over the small étale site of $X$:
$\mathit{DM}_{c}(X,\Lambda)\cong D^{b}_{\mathrm{ctf}}(X,\Lambda)$
(see [CD16, Theorem 6.3.11]).
Example 1.3.
Let $\ell$ be a prime number and define
[…]
where the limit is taken in the $\infty$-category of (stable) $\infty$-categories. In the case where $\ell$ is prime to the residue characteristics of $X$, the rigidity theorem imples that $D^{b}_{c}(X,\mathbf{Z}_{\ell})$ is the usual derived category of constuctible $\ell$-adic sheaves on the small étale site of $X$; see [CD16, Prop. 7.2.21]. The assignment
$X\mapsto D(X,\mathbf{Z}_{\ell})=\mathrm{Ind}(D^{b}_{c}(X,\mathbf{Z}_{\ell}))$
is then a six-functors formalism that satisfies all properties listed in paragraph 1.1. We observe that the dualizable objects are then exactly those complexes whose cohomology objects are what Deligne calls “faisceaux localement constants tordus”. Indeed, the functor
$D^{b}_{c}(X,\mathbf{Z}_{\ell})\to D^{b}_{c}(X,\mathbf{F}_{\ell})\ ,\quad F\mapsto F/\ell$
is conservative (by definition of what $\ell$-adic” means) and symmetric monoidal. This implies that it preserves and detects dualizable objects. Therefore, by virtue of [CD16, Remark 6.3.27], $F$ in $D^{b}_{c}(X,\mathbf{Z}_{\ell})$ is dualizable if and only if $F/\ell$ is locally free with perfect values; for a discussion using condensed techniques, see [BS15, HRS23]. We denote by $D(X,\mathbf{Q}_{\ell})$ the full subcategory of uniquely divisible objects in $D(X,\mathbf{Z}_{\ell})$. It is compactly generated with category of compact objects
$D^{b}(X,\mathbf{Q}_{\ell})=D^{b}_{c}(X,\mathbf{Z}_{\ell})\otimes\mathbf{Q}\,.$
Example 1.4.
Let $\mathcal{U}$ be an ultrafilter on the set of prime numbers. We define for any noetherian $X$:
[…]
(where the colimit and the product are taken in the $\infty$-category of (stable) $\infty$-categories). We define $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ as the full subcategory of $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ by noetherian induction: an object $F$ of $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ belongs to $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ if there is a dense open subscheme $V$ of $U$ such that the restriction $F_{|_{V}}$ is dualizable in $D^{b}_{c}(V,\mathbf{F}_{\mathcal{V}})^{\prime}$ and $F$ is constructible on a complement of $V$. We define finally
$D(X,\mathbf{F}_{\mathcal{U}})=\mathrm{Ind}(D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}}))\,.$
If the set of residue characteristics of $X$ is finite, $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ coincides with the $\mathcal{U}$-indexed ultraproduct of the ordinary categories of constructible sheaves $D^{b}_{c}(X,\mathbf{F}_{\ell})$. If the set of residue characteristics is finite (e.g. is X is of finite type over a semi-local scheme), then $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ coincides with the ultraproduct of the usual categories of étale sheaves $D(X,\mathbf{F}_{\ell})$, where $\ell$ runs over all primes distinct from the residue characteristics of $X$. The corresponding cohomology is then the ultraproduct in the usual sense of étale cohomology groups:
[…]
For an arbitrary scheme $X$, by virtue of [CD16, Cor. 4.5.3], we always have
$\mathit{DM}_{h,lc}(X,\mathbf{F}_{\ell})\cong D^{b}_{c}(X[1/\ell],\mathbf{F}_{\ell})$
where $X[1/\ell]$ is the product of $X$ with the spectrum of $\mathbf{Z}[\frac{1}{\ell}]$. Therefore, we have
[…]
The only non-trivial property of 1.1 that is difficult to check is (i). In the case of quasi-excellent schemes over a semi-local noetherian ring, this follows right away from Orgogozo’s uniform constructibility theorem [Org19, Théorème 3.1.1].
Example 1.5.
If the base field has a complex embedding $k\hookrightarrow\mathbf{C}$, one can define the Betti six-functors formalism
$X\mapsto D_{\textit{Betti}}(X)=D(X(\mathbf{C}),\mathbf{Z})$
by taking derived $\infty$-categories of sheaves of abelian groups on the analytic variety associated to $X$. The associated cohomology is Betti cohomology.
Example 1.6.
If $k$ is of characteristic zero, one can define the de Rham six-functors formalism over schemes of finite type over $k$ through the theory of algebraic $D$-modules. The corresponding cohomology is algebraic de Rham cohomology.
Example 1.7.
If $k$ is a perfect field of characteristic $p>0$, choosing a complete discrete valuation ring of characteristic zero with residue field $k$ (such as the ring of Witt vectors), Berthelot’s theory of arithmetic $D$-modules defines a six-functors formalism for schemes of finite type over $k$. The corresponding cohomology is then Berthelot’s rigid cohomology (i.e. it coincides with Monsky-Washnitzer cohomology for affine smooth $k$-schemes and with crystalline cohomology for smooth and projective ones).
1.8.
If $D$ and $D^{\prime}$ are two six-functors formalisms defined for schemes over $S$, there is a natural notion of morphism of six-functors formalisms from $D$ to $D^{\prime}$; see [DG22]. Informally, this consists of colimit-preserving (hence exact) symmetric monoidal functors
$D(X)\to D^{\prime}(X)$
that are compatible with pullbacks $f^{*}$ for all $f$ as well with direct image with compact support $f_{!}$ for any $f$ separated and of finite presentation. In particular, on noetherian schemes, such morphisms must preserve constructible objets (they preserve dualizable objects, which implies that they preserve constructible objects by noetherian induction). One can prove that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying étale descent [^2] One can see that $\mathit{DM}_{h}$ is initial among $\mathbf{Z}$-linear six-functors formalisms satisfying étale hyperdescent: since we know what is the initial six-functors formalism by virtue of Drew and Gallauer’s [DG22, Theorem 7.14], we see easily that it suffices to prove it for $\mathbf{Q}$-linear coefficients, which follows right away from [CD19, Theorem 16.1.2 and 16.2.18], and for finite coefficients, which follows from the combination of rigidity theorems: the one recalled in Example 1.2 together with those stated in [Ayo14, Bac21]. This implies that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying the properties of 1.1. For $X$ noetherian with residue fields that have a uniform bound on their Galois cohomological dimension, $\mathit{DM}$ and $\mathit{DM}_{h}$ agree anyway.. In particular, for any other six-functor formalism $D$, there is a unique morphism
$r_{D}\colon\mathit{DM}(-)\to D(-)\,.$
We call an object $F$ of $D(X)$ strictly of geometric origin if there is a constructible motivic sheaf $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$ (in particular, such an $F$ is constructible). We will say that $D$ behaves well over $S$ if the functor $r_{D}$ commutes with functors of the form $f_{*}$ for any morphism of finite presentation $f$. All examples of six-functors formalisms given in the present notes are known to behave well: in the case where $D$ is $\mathbf{Q}$-linear and defined on schemes of finite type over a field, this is subsumed by [CD19, Theorem 4.4.25] – see Remark 1.13 below to drop the $\mathbf{Q}$-linearity. For $D=D(-,\Lambda)$ with $\Lambda$ of positive characteristic, this is obvious, by rigidity. This is also true in the case where $D(X)=D(X,\mathbf{Z}_{\ell})$ for any $S$ (see [CD16, Theorem 7.2.11]), or $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ (exercise).
The property of being dualizable is at the core of many constructions and computations. In particular, the fact that any construcible sheaf becomes dualizable over a dense open subscheme is at the center of everything. For objects that are strictly of geometric origin, the choice of such a dualizibility locus can be made motivically, hence, independently of $D$. This is made precise in the following statement.
Proposition 1.9.
Let $D$ be six-functors formalism defined over $S$-schemes. If $F$ is strictly of geometric origin in $D(X)$, then one can find a dense open subscheme $U$ of $X$, independently of $D$, such that the restriction $F_{|U}$ is dualizable in $D(U)$.
Proof.
Let $M$ be constructible in $\mathit{DM}(X)$ such that $F\cong r_{D}(M)$. We choose a dense open immersion $j\colon U\to X$ such that $M_{|U}=j^{*}(M)$ is dualizable in $\mathit{DM}(X)$. Since the functor
$r_{D}\colon\mathit{DM}(U)\to D(U)$
is symmetric monoidal, it preserves dualizable objects, and therefore $F_{|U}\cong r_{D}(M_{|U})$ is dualizable. The open subscheme $U$ clearly depends on $M$ only. ∎
1.10.
The following statement, is a generalization of Deligne’s generic base change theorem for torsion étale sheaves [Del77, Th. Finitude, 1.9]. With this new degree of generality, it may be seen as an independence of $D$ (in particular, independence of $\ell$) result as well; see Corollary 1.14.
Theorem 1.11 (Deligne’s generic base change formula).
Let $f:X\to Y$ be a morphism between separated schemes of finite type over a noetherian base scheme $S$. Let $F$ be a constructible object in $D(X)$. Then there is a dense subscheme $U\subset S$ over which $f_{*}(F)$ becomes constructible and its formation is compatible with any base-change. More precisely, for each $w:S^{\prime}\to S$ factoring through $U$ we have an isomorphism of constructible objects
$v^{*}f_{*}(F)\cong f^{\prime}_{*}u^{*}(F)$
where
[…]
is the associated pull-back diagram.
Proof.
See [Cis21, Theorem 2.4.2] (the proof is stated for $D(X)=\mathit{DM}(X)$ but it only uses the properties listed in paragraph 1.1). The statement in loc. cit. does not say that $f_{*}(F)$ becomes constructible over $U$, but this is what the proof yields anyway. ∎
Remark 1.12.
The proof of the generic base change formula really uses étale descent on constructible sheaves: this implies that, for any quasi-finite morphism $f$. the functors $f_{!}$ and $f_{*}$ are conservative, which is a crucial ingredient in Deligne’s proof. It is not known whether the generic base change formula is true or not for more universal theories of motivic sheaves in which we only have Nisnevich descent, such as Voevodsky’s and Morel’s $\mathit{SH}$.
Remark 1.13.
In fact, a close examination of the proof of the generic base change formula shows that, for a given constructible $F$ in $D(X)$ and a morphism of finite presentation $f:X\to Y$ over some base $S$, one can choose the open subset $U\subset S$ so that the conclusion of the generic base change theorem holds for $f_{*}(F)$ in $D$ and so that, for any morphism of six-functors formalisms $\varphi\colon D\to D^{\prime}$ (with $D^{\prime}$ satisfying the assumption of 1.1 as well), the canonical map
$\varphi(f_{*}(F))\to f_{*}(\varphi(F))$
becomes an isomorphism over $U$ and the conclusion of the generic base change theorem holds for $\varphi(F)$ with respect to $f$ in $D^{\prime}$ (the idea is that we can do the proof of the generic base change formula for $F$ in $D$ and for $\varphi(F)$ in $D^{\prime}$ in parallel and observe that $\varphi$ is compatible with each reduction step, thus giving rise to the same choice of locus $U$ over which the base change formula holds).
Corollary 1.14.
If moreover $F$ is strictly of geometric origin, then the open subscheme $U\subset S$ in the statement of Theorem 1.11 may be chosen indepently of $D$.
Proof.
Let $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$. Apply the generic base change formula to $M$, which gives an open subscheme $U\subset S$ such that the conclusion of Theorem 1.11 holds in $\mathit{DM}$. The conclusion of Theorem 1.11 holds for $r_{D}(M)$ in $D$ as well, by the previous Remark. [^3] For the reader who does not want to verify the claims of Remark 1.13, Corollary 1.14 can be proved as follows in the case where $D$ behaves well: the conclusion of Theorem 1.11 holds for $F=r_{D}(M)$ in $D$ with respect to the motivically chosen $U\subset S$: since $r_{D}\colon\mathit{DM}\to D$ commutes with the six operations, all the base change isomorphisms involving $M$ in $\mathit{DM}$ will yield analogous base change isomorphisms involving $F$ in $D$.
1.15.
We fix a six-functors formalism $D$ defined over $S$-schemes, as in paragraph 1.1. Let $E$ be a vector bundle over $S$ and $\mathbf{P}=\mathbf{P}(E)$ the associated projective bundle over $S$. We denote by $\mathbf{P}^{\wedge}$ the projective bundle associated to the dual of $E$, and we assume given a diagram of $S$-schemes of finite presentation
$X\overset{e}{\to}\bar{X}\overset{\bar{f}}{\to}\mathbf{P}$
(in practice, $e$ will be an open immersion and $\bar{X}$ will be a compactification of $X$ as an $S$-scheme, but, for the moment, $e$ and $\bar{f}$ could be any maps). We let
$h_{\mathit{univ}}\colon\mathbf{H}\hookrightarrow\mathbf{P}\times_{S}\mathbf{P}^{\wedge}$
be the universal hyperplane in $\mathbf{P}$, and write
$\mathbf{P}\overset{\pi}{\leftarrow}\mathbf{P}\times_{S}\mathbf{P}^{\wedge}\overset{\pi^{\wedge}}{\to}\mathbf{P}^{\wedge}$
for the canonical projections. We consider the following diagram, in which all squares are cartesian.
[…]
Observe that both projections $\mathbf{H}\to\mathbf{P}$ and $\mathbf{H}\to\mathbf{P}^{\wedge}$ are smooth. Given a constructible object $F$ in $D(X)$, we then have isomorphisms
$\bar{p}^{*}\,e_{*}(F)\cong\gamma_{*}\,p^{*}(F)$
and
$h^{*}_{\bar{X}}\,\gamma_{*}\,p^{*}(F)\cong\eta_{*}\,h^{*}_{X}\,p^{*}(F)\,.$
The first one is an instance of the smooth base change isomorphism, whereas the second one comes from the following sequence of isomorphisms (using the smoothness of $\mathbf{H}$ over the projective bundle $\mathbf{P}$):
$h^{*}_{\bar{X}}\,\gamma_{*}\,p^{*}(F)$ $\cong h^{*}_{\bar{X}}\,\bar{p}^{*}\,e_{*}(F)$
$\cong(\bar{p}h_{\bar{X}})^{*}\,e_{*}(F)$
$\cong\eta_{*}\,(ph_{X})^{*}(F)$
$\cong\eta_{*}\,h^{*}_{X}\,p^{*}(F)\,.$
This leads to the following statement that goes back to Beilinson [Bei87] and Katz [Kat93].
Theorem 1.16 (shadow of the weak Lefschetz theorem).
Under the assumptions above, there exists a dense open subscheme $U\subset\mathbf{P}^{\wedge}$ with the following property: for any hyperplane $H_{t}$ in $\mathbf{P}$ defined over an $S$-scheme $T$ corresponding to a map $t\colon T\to U$ over $S$, if we form the corresponding diagram of cartesian squares
$X_{T}-H$$X_{T}$$H¯{\bar{X}_{T}-\bar{H}}$$\bar{X}_{T}$$\bar{H}$$\scriptstyle{j}$$\scriptstyle{e_{T}}$$\scriptstyle{u}$$\scriptstyle{i}$$\scriptstyle{\bar{\jmath}}$$\scriptstyle{\bar{\imath}}$
in which $X_{T}$ and $\bar{X}_{T}$ are the pullbacks of $X$ and $\bar{X}$ over $T\to S$ respectively, and $\bar{H}$ is the pullback of $H_{t}$ along the projection $H_{\bar{X}}\to\mathbf{P}^{\wedge}$, we have a canonical isomorphism
$\bar{\jmath}_{!}\,\bar{\jmath}^{*}\,(e_{T})_{*}(F_{T})\cong(e_{T})_{*}\,j_{!}\,j^{*}(F_{T})$
with $F_{T}$ the pullback of $F$ along the projection $X_{T}\to X$. Moreover, if ever $F$ is strictly of geometric origin, the subscheme $U$ can be chosen independently of $D$.
Proof.
We apply the generic base change formula to the restriction of $p^{*}(F)$ on $H_{X}$ and $\eta$, seen over the scheme of hyperplane sections $\mathbf{P}^{\wedge}$. There is thus a dense open subscheme $U\subset\mathbf{P}^{\wedge}$ over which $\eta_{*}\,h^{*}_{X}\,p^{*}(F)$ is constructible and its formation stable under any base change along any map $T\to U\subset\mathbf{P}^{\wedge}$. Applying the generic base change formula to $p^{*}(F)$ with respect to the map $\gamma$ over $\mathbf{P}^{\wedge}$, we may shrink $U$ so that the formation of $\gamma_{*}\,p^{*}(F)$ is stable under any base change along any map $T\to U\subset\mathbf{P}^{\wedge}$ as well.
Let us consider an $S$-scheme $T$ together with a map $t\colon T\to U$ defining an hyperplane $H_{t}$ in $\mathbf{P}$. We obtain pullback squares of the form
[…]
inducing base change isomorphisms
$t^{*}_{\bar{X}}\,\gamma_{*}\,p^{*}(F)\cong(e_{T})_{*}\,\bar{\imath}^{*}\,p^{*}(F)=(e_{T})_{*}(F_{T})$
and
$t^{*}_{\bar{H}}\,\eta_{*}\,h^{*}_{X}\,p^{*}(F)\cong u_{*}\,t^{*}_{H}\,p^{*}(F)=u_{*}\,i^{*}(F_{T})\,.$
We claim that these induce a canonical isomorphism:
$\bar{\imath}_{*}\,\bar{\imath}^{*}\,(e_{T})_{*}(F_{T})\cong(e_{T})_{*}\,i_{*}\,i^{*}(F_{T})\,.$
Indeed, we have:
$\bar{\imath}_{*}\,\bar{\imath}^{*}\,(e_{T})_{*}(F_{T})$ $\cong\bar{\imath}_{*}\,\bar{\imath}^{*}\,t^{*}_{\bar{X}}\,\gamma_{*}\,p^{*}(F)$
$\cong\bar{\imath}_{*}\,t^{*}_{\bar{H}}\,h^{*}_{\bar{X}}\,\gamma_{*}\,p^{*}(F)$
$\cong\bar{\imath}_{*}\,t^{*}_{\bar{H}}\,\eta_{*}\,h^{*}_{X}\,p^{*}(F)$
$\cong\bar{\imath}_{*}\,u_{*}\,i^{*}(F_{T})$
$\cong(e_{T})_{*}\,i_{*}\,i^{*}(F_{T})\,.$
On the other hand, he have a canonical morphism of cofiber sequences:
$\bar{\jmath}_{!}\,\bar{\jmath}^{*}(e_{T})_{*}(F_{T})$$(e_{T})_{*}(F_{T})$$\bar{\imath}_{*}\,\bar{\imath}^{*}\,(e_{T})_{*}(F_{T})$$(e_{T})_{*}\,j_{!}\,j^{*}(F_{T})$$(e_{T})_{*}(F_{T})$$(e_{T})_{*}\,i_{*}\,i^{*}(F_{T})$$\scriptstyle{\cong}$
in which the vertical map on the right hand side is an isomorphism by the computation we just did. Since the middle vertical map is the identity, this proves that the vertical map on the left hand side is an isomorphism. The fact the open subscheme $U\subset\mathbf{P}^{\wedge}$ may be chosen independently of $D$ if ever the sheaf $F$ is strictly of geometric origin follows from Corollary 1.14. ∎
Corollary 1.17.
Let $k$ be an infinite field, and $X$ a quasi-projective scheme over $k$. We choose an open immersion $e:X\to\bar{X}$ into a projective $k$-scheme, and a constructible motivic sheaf $M$ in $\mathit{DM}(X)$. There exists a dense open subset $U$ (hence with at least one rational point) of the family of hyperplane sections of $\bar{X}$, such that, for any $\bar{H}$ in $U$, with $H=\bar{H}\cap X$ and with corresponding open complements $\bar{\jmath}\colon\bar{X}-\bar{H}\to\bar{X}$ and $j\colon X-H\to X$, and for any six functors formalism $D$ as in paragraph 1.1, if we put $F=r_{D}(M)$ for the realization of $M$ in $D$, then there is a canonical isomorphism
$\bar{\jmath}_{!}\,\bar{\jmath}^{*}\,e_{*}(F)\cong e_{*}\,j_{!}\,j^{*}(F)\,.$
In particular, the cohomology with compact support of $\bar{X}-\bar{H}$ with coefficients in $e_{*}(F)$ coincides with the cohomology of $X$ with coefficients in $j_{!}\,j^{*}(F)$.
1.18.
A fundamental situation that will come back eventually in these notes is the following one. Let $F$ be in $D(X)$ for an affine scheme of finite type $X$ over a field $k$. We assume chosen an affine open embedding $e\colon X\to\bar{X}$ into a projective scheme. A hyperplane section $\bar{H}$ of $\bar{X}$ with trace $H=X\cap\bar{H}$ is said to be transverse to $F$ if by forming the corresponding pullback squares below
$X¯H¯{\bar{X}-\bar{H}}X¯{\bar{X}}H¯{\bar{H}}$$\scriptstyle{j}$$\scriptstyle{e}$$\scriptstyle{u}$$\scriptstyle{i}$$\scriptstyle{\bar{\jmath}}$$\scriptstyle{\bar{\imath}}$
there is a canonical isomorphism
$\bar{\jmath}_{!}\,\bar{\jmath}^{*}\,e_{*}(F)\cong e_{*}\,j_{!}\,j^{*}(F)\,.$
Theorem 1.16 tells that there is a dense open subspace of such hyperplane sections and therefore that, possibly after base change along a finite extension of $k$, one can provide such nice hyperplane sections. Furthemore, if ever $F$ is strictly of geometric origin, one can choose a dense open subspace of hyperplane sections in good position with respect to $F$ independently of $D$.
2. Perverse sheaves
2.1.
Let $D$ be a six-functors formalism defined on schemes of finite presentation over a base $S$. A $t$-structure on $D$ is a family of $t$-structures on $D(X)$ for each $X$ with the following properties (we adopt cohomological conventions):
  • (a)
    for any scheme $X$ any integer $n\in\mathbf{Z}$, the Tate object $\mathbf{1}_{X}(n)$ is in the heart of $D(X)$;
  • (b)
    for any scheme $X$, the tensor product functor preserves objects concentrated in degree $\leq 0$ in the sense of the $t$-structure (i.e. tensoring with an object of the heart is right $t$-exact in $D(X)$);
  • (c)
    for any map $f\colon X\to Y$, the functor $f^{*}\colon D(Y)\to D(X)$ is $t$-exact;
  • (d)
    for any separated morphism of finite presentation $f:X\to Y$, the functor $f_{!}\colon D(X)\to D(Y)$ is right $t$-exact;
  • (e)
    the truncation functors preserve dualizable objects and all dualizable objects of $D(X)$ are bounded for the $t$-structure;
  • (f)
    producing the dual of dualizable objects $M\mapsto M^{\wedge}$ is a $t$-exact functor.
As a consequence, Tate twists $M\mapsto M(n)$ are $t$-exact functors. Any constructible object $F$ of $D(X)$ is bounded. Indeed, if we define $f$ as the coproduct of the embeddings of the strata over which $F$ becomes dualizable, then $f^{*}(F)$ is dualizable in $D(X^{\prime})$, and since $f^{*}$ is conservative (by localization) and $t$-exact, this proves the claim. Remark that condition (f) is only reasonnable with $\mathbf{F}$-linear coefficients for some field $\mathbf{F}$. [^4] Condition (f) can be avoided but we would need two $t$-structures whose restrictions to dualizable objects would be exchanged through ordinary duality, and that would only make notations worse. We denote by $D^{\leq i}_{c}(X)$ (by $D^{\geq i}_{c}(X)$ respectively) the full subcategory of $D_{c}(X)$ that consists of constructible objects with cohomology concentrated in degrees $\leq i$ (in degrees $\geq i$, respectively) with respect to the $t$-structure above. Such a $t$-structure will be called ordinary.
Example 2.2.
It is well known that, to say the least, the existence of such a $t$-structure is very conjectural as far as $\mathit{DM}$ is concerned. However, $D(X)=D(X,\mathbf{Q}_{\ell})$, $D(X)=D(X,\mathbf{F}_{\ell})$ and $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ have canonical $t$-structures coming from the usual $t$-structures on the derived category of sheaves of $\Lambda$-modules on the small étale site of $X$, with $\Lambda$ and finite commutative ring. If moreover $X$ is noetherian, all the properties listed above will be fulfilled in these cases.
2.3.
Let $k$ be a field and $D$ a six-functors formalism as in 1.1 defined on schemes of finite type over $k$ and equipped with a $t$-structure in the sense of 2.1. We can then reproduce word for word section 2.2 of [BBDG18]. We will always consider here the autodual perversity. The perverse $t$-structure on $D(X)$ is characterized by noetherian induction, using the following facts:
  • (i)
    The perverse $t$-structure coincides with the usual one on $D(X)$ for any $k$-scheme $X$ of dimension $\leq 0$;
  • (ii)
    For any closed immersion $i\colon Z\to X$, the functor $i_{!}\colon D(Z)\to D(X)$ is $t$-exact with respect to the perverse $t$-structure, and for any open immersion $j\colon U\to X$, the functor $j^{*}\colon D(X)\to D(U)$ is $t$-exact with respect to the perverse $t$-structure;
  • (iii)
    For any regular scheme $X$ of pure dimension $d$, any dualizable object $F$ of $D(X)$ that is also in the heart of the usual $t$-strucure defines an object $F[-\dim(X)]$ in the heart of the perverse $t$-structure of $D_{c}(X)$.
More precisely, we define the perverse $t$-structure on $D(X)$ by induction on the dimension of $X$ (together with properties (i), (ii) and (iii) above). For $X$ of dimension $\leq 0$, this is the ordinary $t$-structure (the one from 2.1). For $X$ of dimension $d>0$, assuming that the perverse $t$-structure is defined on $D(Z)$ for any scheme $Z$ of dimension <d<d, we define the perverse $t$-structure on $D(X)$ as follows. For any nowhere dense closed subscheme $i\colon Z\hookrightarrow X$ with open complement $j\colon U\hookrightarrow X$ such that $U_{\mathit{red}}$ is regular, we denote by $D_{U}(X)$ the full subcategory of $D(X)$ spanned by constructible objects $F$ with the property that $j^{*}(F)$ is dualizable. If $D^{\wedge}(U)$ denotes the full subcategory of dualizable objects of $D(U)$, the functors $i_{*}$ and $j^{*}$ expressing $D(X)$ as a recollement of $D(U)$ and $D(Z)$ in the sense of [BBDG18, 1.4.3] exhibit $D_{U}(X)$ as a recollement of $D^{\wedge}(U)$ and $D(Z)$. If $U=\coprod_{i}U_{i}$ is the decomposition of $U$ into its irreducible components, then
$D^{\wedge}(U)\cong\prod_{i}D^{\wedge}(U_{i})$
and we define the perverse $t$-structure on $D^{\wedge}(U)$ as the product perverse $t$-structure, where the perverse $t$-structure on each $D^{\wedge}(U_{i})$ is defined by the property that its heart consists of those objects of the form $F[-\dim(U_{i})]$ with $F$ in the heart of the ordinary $t$-structure on $D(U_{i})$. We have a perverse $t$-structure on $D(Z)$ by induction, and thus a perverse $t$-structure on $D_{U}(X)$, by [BBDG18, Théorème 1.4.10]. If $Z^{\prime}$ is a nowhere dense open subscheme of $X$ containing $Z$ and if $U^{\prime}$ is the complement of $Z^{\prime}$, since the extension by zero functor $D(Z)\to D(Z^{\prime})$ is $t$-exact (by induction on dimension, property (ii) holds true), and since the restriction functor $D^{\wedge}(U)\to D^{\wedge}(U^{\prime})$ is obviously $t$-exact, we see that the inclusion functor
$D_{U}(X)\hookrightarrow D_{U^{\prime}}(X)$
is $t$-exact for the perverse $t$-structures. Writing
[…]
where the filtered colimit is indexed by dense open subschemes $U\subset X$ such that $U_{\mathit{red}}$ is regular thus defines a $t$-structure on $D_{c}(X)$. It is clear that properties (i), (ii) and (iii) are true by construction.
We denote by $^{p}D^{\leq i}_{c}(X)$ (by $^{p}D^{\geq i}_{c}(X)$ respectively) the full subcategory of $D_{c}(X)$ that consists of constructible objects with cohomology concentrated in degrees $\leq i$ (in degrees $\geq i$, respectively) with respect to the perverse $t$-structure defined above.
An object of $D(X)$ is called perverse if it is in the heart of the perverse $t$-structure on $D_{c}(X)$.
Lemma 2.4.
For any closed immersion $i\colon Z\to X$, the functor $i^{*}\colon D(X)\to D(Z)$ is right $t$-exact with respect to the perverse $t$-structure.
Proof.
This is obvious since its right adjoint $i_{*}$ is (left) $t$-exact with respect to the perverse $t$-structure. ∎
2.5.
Observe that, by generic base change, the functors $f_{*}$ preserve constructible objects for any morphism $f$ between $k$-schemes of finite type. This implies that, for any separated morphism of finite type $f\colon X\to Y$, the functor $f^{!}$ preserves constructible objects in $D$: since this is a Zariski-local property, it suffices to check this property for $f$ quasi-projective. By relative purity, for $f$ smooth, we have $f^{!}(\mathbf{1})\otimes f^{*}\cong f^{!}$ with the object $f^{!}(\mathbf{1})$ $\otimes$-invertible (hence dualizable and therefore constructible), so that it suffices to check that $i^{!}$ preserves constructible objects for $i\colon Z\to X$ a closed immersion. If $j\colon U\to X$ is the open complement, the fiber sequence
$i^{!}\to i^{*}\to i^{*}\,j_{*}\,j^{*}$
shows that $i^{!}$ preserves constructible objects. We also see that the internal Hom preserve constructible objects: For $F$ and $G$ constructible in $D(X)$, we choose an open immersion $j\colon U\to X$ so that $j^{*}(F)$ is dualizable. The internal Hom $\mathit{Hom}(F,G)$ then fits into a fiber sequence
$i_{*}\,i^{!}\mathit{Hom}(F,G)\to\mathit{Hom}(F,G)\to j_{*}\,j^{*}\mathit{Hom}(F,G)$
that is isomorphic to
$i_{*}\mathit{Hom}(i^{*}(F),i^{!}(G))\to\mathit{Hom}(F,G)\to j_{*}\mathit{Hom}(j^{*}(F),j^{*}(G))\,.$
We see that $\mathit{Hom}(i^{*}(F),i^{!}(G))$ by induction on the Krull dimension whereas the object $\mathit{Hom}(j^{*}(F),j^{*}(G))$ is the tensor product of the dual of the dualizable object $j^{*}(F)$ with the constructible object $j^{*}(G)$, hence is constructible. We can thus define Verdier duality in $D$ as the functor
$D_{X}\colon D_{c}(X)^{\mathit{op}}\to D_{c}(X)\ ,\quad F\mapsto\mathit{Hom}(F,a^{!}(\mathbf{1}))$
where $a\colon X\to\mathrm{Spec}(k)$ is the structural map. Observe that, since
$f^{!}\mathit{Hom}(A,B)\cong\mathit{Hom}(f^{*}(A),f^{!}(B))$
and
$\mathit{Hom}(f_{!}(A),B)\cong f_{*}\mathit{Hom}(A,f^{!}(B))\,$,
we see that Verdier duality commutes with $f_{*}$ for $f$ proper.
Remark 2.6.
On regular schemes, Verdier duality preserves dualizable objects: For $F$ dualizable in $D(X)$ with $X_{\mathit{red}}$ regular, we have $D_{X}(\mathbf{1}_{X})=\mathbf{1}_{X}(d)[2d]$ with $d=\dim X$ (it suffices to prove this for $D=DM$ since $r_{D}\colon DM\to D$ preserves the six operations) so that
$D_{X}(F)\cong F^{\wedge}\otimes\mathbf{1}_{X}(d)[2d]=F^{\wedge}(d)[2d]$
with $F^{\wedge}$ the ordinary dual of $F$. In particular, we have
$D_{X}(F[d]])\cong D_{X}(F)[-d]\cong F^{\wedge}[d](d)\,.$
Definition 2.7.
An Artin-Verdier six-functors formalism is a six-functors formalism $D$ defined for schemes of finite type over a field $k$ as in 1.1 equipped with an ordinary $t$-structure in the sense of 2.1 with the following two extra properties:
  • (i)
    Verdier duality holds in $D$: for any scheme of finite type $X$ over $k$, Verdier duality induces an equivalence
    $D_{X}\colon D_{c}(X)^{\mathit{op}}\cong D_{c}(X)\,.$
  • (ii)
    Artin vanishing holds in $D$: for any affine morphism of $k$-schemes of finite type $f\colon X\to Y$ the induced functor $f_{*}\colon D_{c}(X)\to D_{c}(Y)$ is right $t$-exact with respect to the perverse $t$-structure.
Example 2.8.
$D(-,\mathbf{F})$ with $\mathbf{F}$ a finite field, $D(-,\mathbf{Q}_{\ell})$ with $\ell$ any prime number, and $D(-,\mathbf{F}_{\mathcal{U}})$ with $\mathcal{U}$ any ultrafilter on the set of prime numbers, as defined in 1.3 and 1.4 are Artin-Verdier six-functors formalisms – properties (i) and (ii) are even known to hold over any quasi-excellent base, after the work of Gabber [ILO14, Exposé XVII, Théorème 0.2 and Exposé XV, Théorème 1.1.2]. Other examples are the de Rham six-functors formalism, the $\mathbf{F}$-linear version of Betti six-functors formalism (for any field of coefficient $\mathbf{F}$), and the rigid six-functors formalism.
2.9.
From now on, we fix an Artin-Verdier six-functors formalism $D$. We consider the associated perverse $t$-structure on $D$ (see 2.3).
Remark 2.10.
We have formulas of the form
$f^{!}\,D_{X}\cong D_{Y}\,f^{*}\quad\text{and}\quad D_{Y}\,f_{!}\cong f_{*}\,D_{X}$
for any morphism of $k$-schemes of finite type $f\colon X\to Y$. Since Verdier duality is an equivalence, this implies dually that
$f^{*}\,D_{X}\cong D_{Y}\,f^{!}\quad\text{and}\quad D_{Y}\,f_{*}\cong f_{!}\,D_{X}$
holds as well. Given an immersion $\imath\colon S\to X$ with $S_{\mathit{red}}$ regular, and $F$ constructible in $D(X)$ such that $\imath^{*}(F)$ (with $\imath^{!}(F)$, respectively) is dualizable, we observe that
$D_{S}\,\imath^{*}(F)\cong\imath^{!}\,D_{S}(F)\quad\text{($D_{S}\,\imath^{!}(F)\cong\imath^{*}\,D_{S}(F)$, respectively)}$
is dualizable as well: it is the tensor product of the $\otimes$-invertible object $D_{S}(\mathbf{1}_{S})\cong\mathbf{1}_{S}(\dim S)[2\dim S]$ with the ordinary dual of $i^{*}(F)$ (of $i^{!}(F)$, respectively); see 2.6.
Proposition 2.11.
Verdier duality is $t$-exact with respect to the perverse $t$-structure.
Proof.
We proceed by induction on the Krull dimension $d$ of $X$. For $d\leq 0$, all constructible objects are dualizable and Verdier duality coincides with ordinary duality, and the bidual of a dualizable object $F$ is canonically isomorphic to $F$. Let us assume that $d>0$. We choose a nowhere dense closed subscheme $i\colon Z\to X$ with open complement $j\colon U\to X$ with $U_{\mathit{red}}$ regular. We consider the category $D_{U}(X)$ of constructible objects $F$ in $D(X)$ such that $j^{*}(F)$ is dualizable. We observe that Verdier duality induces a $t$-exact functor with respect to the perverse $t$-structure
$D_{c}(Z)^{\mathit{op}}\to D_{c}(Z)$
by induction on the Krull dimension, and an equivalence
$D_{U}(U)^{\mathit{op}}\cong D_{U}(U)$
because, up to a Tate twist, Verdier duality on $D_{U}(U)$ simply is ordinary duality restricted to dualizable objects; see Remark 2.6. In fact, the perverse $t$-structure on $D_{U}(U)$ is the ordinary $t$-structure shifted by $\dim U$: the $t$-structure whose heart is
$D^{\leq\dim U}_{U}(U)\cap D^{\geq\dim U}_{U}(U)\,.$
Verdier duality induces an equivalence $D_{U}(U)^{\mathit{op}}\cong D_{U}(U)$ that is $t$-exact with respect to the perverse $t$-structure because of formula $D_{X}(F[d]])\cong F^{\wedge}[d](d)$ for any $F$ in the ordinary heart of $D_{U}(U)$. Verdier duality is $t$-exact with respect to the perverse $t$-structure on $D(Z)$ by induction on $d$. Verdier duality restricts to $D_{U}(X)$, and since the inclusion $D_{U}(X)\hookrightarrow D(X)$ is $t$-exact with respect to the perverse $t$-structure by construction, it suffices to prove that Verdier duality is $t$-exact on $D_{U}(X)$. In other words, we want to prove that an object $F$ of $D_{U}(X)$ is in $^{p}D^{\leq 0}_{U}(X)$ if and only if is Verdier dual $D_{X}(F)$ is in $^{p}D^{\geq 0}_{U}(X)$. By definition of glued $t$-structures, an object $F$ is in $^{p}D^{\leq 0}_{U}(X)$ if only if $i^{*}(F)$ lies in $^{p}D^{\leq 0}(Z)$ and $j^{*}(F)$ lies in $^{p}D^{\leq 0}_{U}(U)$. Dually, an object $F$ is in $^{p}D^{\leq 0}_{U}(X)$ if only if $i^{!}(F)$ lies in $^{p}D^{\geq 0}(Z)$ and $j^{*}(F)$ lies in $^{p}D^{\leq 0}_{U}(U)$. Since Verdier duality exchanges $i^{*}$ and $i^{!}$ and commutes with $j^{*}$, this achieves the proof. ∎
Corollary 2.12.
For any affine morphism of $k$-schemes of finite type $f\colon X\to Y$ the induced functor $f_{!}\colon D_{c}(X)\to D_{c}(Y)$ is left $t$-exact with respect to the perverse $t$-structure.
Corollary 2.13.
For any finite morphism $f\colon X\to Y$, the functor $f_{!}=f_{*}\colon D(X)\to D(Y)$ is $t$-exact with respect to the perverse $t$-structure.
Lemma 2.14.
Let $X$ be of pure dimension $d$, with $X_{\mathit{red}}$ regular and $F$ dualizable in $D(X)$. Then $F$ is perverse (is in $^{p}D^{\leq 0}(X)$) if and only if F[d]F[-d] is in the heart of the ordinary $t$-structure (is in $D^{\leq 0}(X)$, respectively).
Proof.
We will prove the respective case first. Let us assume that $F$ is in $^{p}D^{\leq 0}(X)$ but not in $^{p}D^{<0}(X)$. There is smallest integer $n$ such that $F$ lies in $D^{\leq n}_{c}(X)$. Then F[n]F[-n] is in $D^{\leq 0}_{c}(X)$ and thus F[dn]F[d-n] is in $^{p}D^{\leq 0}(X)$, which yields $n\leq d$.
Let us assume now that $F$ is perverse. Since $F$ is dualizable and $X_{\mathit{red}}$ regular of dimension $d$, we have $D_{X}(F)\cong F^{\wedge}(d)[2d]$ (Remark 2.6). By virtue of Proposition 2.11, this means that
$F^{\wedge}[2d]=(F[-d])^{\wedge}[d]$
is dualizable and perverse as well. This means that both F[d]F[-d] and $(F[-d])^{\wedge}$ are dualizable and in $D^{\leq 0}(X)$. Since ordinary duality of dualizable objects is $t$-exact with respect to the ordinary $t$-structure, this implies that F[d]F[-d] lies in the heart of the ordinary $t$-structure. ∎
Proposition 2.15.
Let $F$ be a constructible object of $D(X)$ and $n\in\mathbf{Z}$ an integer. The following conditions are equivalent.
  • (i)
    $F$ belongs to $^{p}D^{\leq 0}_{c}(X)$ (to $^{p}D^{\geq 0}_{c}(X)$, respectively);
  • (ii)
    For any irreducible closed subscheme $T\subset X$, there is a dense open subscheme $S$ in $T$ inducing an embedding $\imath_{S}\colon S\to X$ such that $\imath^{*}_{S}(F)$ belongs to $D^{\leq\dim S}_{c}(S)$ (such that $\imath^{!}_{S}(F)$ belongs to $D^{\geq\dim S}_{c}(S)$, respectively).
  • (iii)
    For any finite stratification $\mathcal{S}$ of $X$ by strata that are irreducible with regular reduction, and for any embedding $\imath_{S}\colon S\to X$ with $S\in\mathcal{S}$, assuming that both $i^{*}_{S}(F)$ and $i^{!}_{S}(F)$ are dualizable, we have $\imath^{*}_{S}(F)$ in $D^{\leq\dim S}_{c}(S)$ (we have $\imath^{!}_{S}(F)$ in $D^{\geq\dim S}_{c}(S)$, respectively).
  • (iv)
    There exists a finite stratification $\mathcal{S}$ of $X$ by strata that are irreducible with regular reduction, such that, for any embedding $\imath_{S}\colon S\to X$ with $S\in\mathcal{S}$, both $i^{*}_{S}(F)$ and $i^{!}_{S}(F)$ are dualizable and we have $\imath^{*}_{S}(F)$ in $D^{\leq\dim S}_{c}(S)$ (we have $\imath^{!}_{S}(F)$ in $D^{\geq\dim S}_{c}(S)$, respectively).
Proof.
Let $Y$ be a scheme of finite type over $k$. For any constructible object $K$ in $D(Y)$, there is a dense open subscheme $U$ of $Y$ such that the restriction of $K$ on $U$ is dualizable in $D(U)$. Since any scheme of finite over scheme is excellent, there is a dense subscheme $U$ of $Y$ that is a finite disjoint uinion of irreducible schemes whose reductions are regular. Therefore, there exists at least one finite stratification $\mathcal{S}$ of $X$ by strata that are irreducible with regular reduction, such that, for any embedding $\imath_{S}\colon S\to X$ with $S\in\mathcal{S}$, both $i^{*}_{S}(F)$ and $i^{!}(S)$ are dualizable. The fact that we have equivalences (i)–(iv) in the case of $^{p}D^{\leq 0}_{c}(X)$ follows right away from Lemmata 2.4 and 2.14. Since Verdier duality exchanges $^{p}D^{\leq 0}_{c}(X)$ and $^{p}D^{\geq 0}_{c}(X)$ (because it is $t$-exact), the respective case follows now from the last part of Remark 2.10. ∎
Corollary 2.16.
We have the following exactness properties with respect to the perverse $t$-structure.
  • (a)
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f^{*}\colon D_{c}(Y)\to D_{c}(X)$ is right $t$-exact and the exceptional pullback functor $f^{!}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
  • (b)
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f_{!}\colon D_{c}(Y)\to D_{c}(X)$ is right $t$-exact and $f_{*}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
Proof.
To prove (a), we use criterium (ii) of Proposition 2.15. Let $T$ be a closed subscheme of $X$. There is a dense open subscheme $U$ of the closure of the image of $T$ in $Y$ such that $\imath_{U}^{*}(F)$ belongs to $D^{\leq\dim U}_{c}(U)$. Therefore, if $S$ denotes the pullback of $U$ in $T$, we have $\imath_{U}^{*}\,f^{*}(F)$ in $D^{\leq\dim U}_{c}(S)$. Since $\dim S=\dim U$ (because $f$ is quasi-finite and $S$ dominates $U$ by construction), this proves that $f^{*}(F)$ is in $^{p}D^{\leq 0}(X)$. By Verdier duality, this proves that $f^{!}$ is left $t$-exact. Property (b) follows straight away from property (a). ∎
Corollary 2.17.
If $f\colon X\to Y$ is étale, then $f^{*}$ is $t$-exact with respect to the perverse $t$-structure.
Corollary 2.18.
If $f\colon X\to Y$ is both affine and quasi-finite, then both functors $f_{!}$ and $f_{*}$ are $t$-exact with respect to the perverse $t$-structure.
Proposition 2.19.
If $f\colon X\to Y$ is a finite flat morphism with both $X_{\mathit{red}}$ and $Y_{\mathit{red}}$ regular, then, for any $F$ that is both perverse and dualizable in $D(Y)$, there is an isomorphism $f^{!}(F)\cong f^{*}(F)$ and $f^{*}(F)$ is both perverse and dualizable in $D(X)$.
Proof.
There is always a canonical morphism
$f^{!}(\mathbf{1}_{Y})\otimes f^{*}(F)\to f^{!}(F)\,.$
The latter is an isomorphism for $F$ dualizable (this is a nice exercise that does not use any property of $f$) and $f^{!}(\mathbf{1}_{Y})\cong\mathbf{1}_{X}$ by relative purity (this times using that both $X$ and $Y$ are regular up to reduction and of same dimension). If $F$ is dualizable in $D(Y)$, since $Y_{\mathit{red}}$ is regular, it is perverse if and only if $F[-\dim Y]$ is in the heart of the ordinary $t$-structure (Lemma 2.14). The functor $f^{*}$ preserves dualizable because it is symmetric monoidal. Since $f^{*}$ is $t$-exact for the ordinary $t$-structure, this proves the proposition. ∎
Proposition 2.20.
Let $a\colon X\to Y$ be a flat affine morphism between $k$-schemes of finite type, of relative dimension $d$.
  1. (a)
    For any $F$ in $D^{\leq 0}_{c}(X)$, there is a dense open subscheme $V$ of $Y$ such that the restriction $a_{*}(F)_{|_{V}}$ belongs to $D^{\leq d}_{c}(V)$.
  2. (b)
    For any $F$ in $D^{\geq 0}_{c}(X)$, such that $F[\dim X]$ belongs to $^{p}D^{\geq 0}_{c}(X)$, there is a dense open subscheme $V$ of $Y$ such that the restriction $a_{!}(F)_{|_{V}}$ belongs to $D^{\geq d}_{c}(V)$.
Moreover, in both cases, if $F$ is strictly of geometric origin, then the open subscheme $V$ may be chosen independently of $D$.
Proof.
For (a), We proceed by induction on $d$. Since we are allowed to shrink $Y$ at will, we may assume $Y_{\mathit{red}}$ to be regular. Since $D$ is insensitive to nil-immersions, we may as well assume that $Y$ is regular. For $d\leq 0$, this is obvious: since $X$ is affine and quasi-finite over $Y$, we know that $a_{*}$ is $t$-exact by Corollary 2.18. For $d>0$, given $F$ constructible in $D^{\leq 0}(X)$, we choose $i\colon Z\to X$ a nowhere dense closed immersion with affine complement $j\colon U\to X$ such that $U_{\mathit{red}}$ is regular and everywhere of dimension $\dim X$, and $j^{*}F$ is dualizable in $D(U)$ (if $F$ is strictly of geometric origin we make this choice in $\mathit{DM}$ to make this independent of $D$, by Proposition 1.9). We then have $j^{*}F[\dim U]$ in $^{p}D^{\leq 0}(U)$ hence $j_{!}\,j^{*}F[\dim U]$ is in $^{p}D^{\leq 0}(X)$. Therefore, Artin vanishing tells us that $a_{*}\,j_{!}\,j^{*}F[\dim U]$ belongs to $^{p}D^{\leq 0}(Y)$. Shrinking $Y$, we may impose that $a_{*}\,j_{!}\,j^{*}F$ is dualizable (if $F$ is strictly of geometric origin, so is $a_{*}\,j_{!}\,j^{*}F$ and we make this shrinking at the motivic level, as in the proof of 1.9). Since $d=\dim U-\dim Y$, by virtue of Lemma 2.14, this means that $a_{*}\,j_{!}\,j^{*}F[d]$ belongs to $D^{\leq 0}(Y)$, hence $a_{*}\,j_{!}\,j^{*}F$ to $D^{\leq d}(Y)$. By virtue of Grothendieck’s generic flatness theorem, shrinking $Y$, we may assume that $Z$ is flat over $Y$ as well (this choice is purely geometric and does not depend on $F$ nor on $D$). By induction, since the relative dimension of $Z$ over $Y$ is <d<d, possibly after shrinking $Y$ one more time, we have $(ai)_{*}\,i^{*}(F)$ in $D^{\leq d-1}(Y)$. The cofiber sequence
$a_{*}\,j_{!}\,j^{*}F\to a_{*}F\to a_{*}\,i_{*}\,i^{*}F$
thus implies that $a_{*}(F)$ belongs to $D^{\leq d}(Y)$.
In order to prove (b), this is more direct. Let $F$ be in $D^{\geq 0}_{c}(X)$, such that $F[\dim X]$ belongs to $^{p}D^{\geq 0}_{c}(X)$. Since $a_{!}$ is left $t$-exact with respect to the perverse $t$-structure (Cor. 2.12), we have $a_{!}(F)[\dim X]$ in $^{p}D^{\geq 0}_{c}(Y)$. Shrinking $Y$, we may assume that $a_{!}(F)$ is dualizable and that $Y_{\mathit{red}}$ is regular, which implies by Lemma 2.14 that
$a_{!}(F)[\dim X-\dim Y]=a_{!}(F)[d]$
belongs to $D^{\geq 0}_{c}(Y)$. ∎
Remark 2.21.
Let $a\colon X\to\mathrm{Spec}(k)$ be a separated morphism of finite type. For any $F$ in $D(X)$, we can define $\Gamma(X,F)=a_{*}(F)$ and $\Gamma_{c}(X,F)=a_{!}(F)$. Taking cohomology in the sense of the ordinary $t$-structure, we get objects
$H^{i}(X,F)=H^{i}(\Gamma(X,f))\quad\text{and}\quad H^{i}_{c}(X,F)=H^{i}(\Gamma_{c}(X,F))$
in the heart $D^{\heartsuit}(\mathrm{Spec}(k))=D^{\leq 0}(\mathrm{Spec}(k))\cap D^{\geq 0}(\mathrm{Spec}(k))$.
The following statement was observed by Beilinson [Bei87]. Recall the notion of hyperplane section transverse to an object of $D$, as defined in 1.18.
Theorem 2.22.
Let $X$ be an affine scheme of finite type over $k$ and $e\colon X\to\bar{X}$ an affine open immersion into a projective scheme. We consider given an affine open immersion $u\colon U\to X$ with $U_{\mathit{red}}$ regular. We assume that $F$ is dualizable and in the heart of the ordinary $t$-structure on $D_{c}(U)$. If $j\colon X-H\hookrightarrow X$ denotes the embedding of the complement of a hyperplane section that is transverse to $u_{!}F$, then $H^{q}(X,j_{!}\,j^{*}u_{!}F)=0$ for any $q\neq\dim X$.
Proof.
Let $d=\dim X$. Keeping track of the notations of 1.18, the transversality assumption on $H$ implies that we have canonical isomorphisms
$H^{q}_{c}(\bar{X}-\bar{H},e_{*}\,u_{!}F)\cong H^{q}(X,j_{!}\,j^{*}\,u_{!}F)$
with both $X$ and $\bar{X}-\bar{H}$ affine of dimension $d$. By virtue of Proposition 2.20 (a), since all functors $_{!}, jj^{*}$ and $u_{!}$ are $t$-exact with respect to the ordinary $t$-structure, $H^{q}(X,j_{!}\,j^{*}\,u_{!}F)$ vanishes for $q>d$. On the other hand, by virtue of Lemma 2.14, we see that $F[d]$ is perverse on $U$, and since both $u$ and $e$ are affine immersions, Corollaries 2.17 and 2.18 imply that $e_{*}\,u_{!}F[d]$ is perverse on $\bar{X}-\bar{H}$. Therefore, Proposition 2.20 (b) implies that $H^{q}_{c}(\bar{X}-\bar{H},e_{*}u_{!}F)$ vanishes for $q<d$. ∎
3. Uniformly perverse motivic generators
3.1.
Let $X$ be a scheme of finite type over a field $k$. We denote by $M(X)$ the motive of $X$, i.e. the object of $\mathit{DM}(\mathrm{Spec}(k))$ defined by
$M(X)=a_{!}\,a^{!}(\mathbf{Z})$
with $a\colon X\to\mathrm{Spec}(k)$ the structural morphism. [^5] If we define $\mathit{DM}(X)$ through $h$-motives, its is a $(\mathbf{P}^{1},\infty)$-stabilization of the $\mathbf{A}^{1}$-localization of the derived $\infty$-category of sheaves of abelian groups on the big site of $k$ with respect to the $h$-topology. The object $M(X)$ is then what remains of the sheaf represented by $X$ through all this process. This construction defines a functor from the category of schemes of finite type over $k$ to the stable $\infty$-category $\mathit{DM}(\mathrm{Spec}(k))$. Given a closed subscheme $Z$ of $X$, we define $M(X,Z)$ through the cofiber sequence
$M(Z)\to M(X)\to M(X,Z)$
For any object $A$ in $\mathit{DM}(\mathrm{Spec}(k))$, we thus have a fiber sequence in the derived category of abelian groups
$\mathrm{RHom}(M(X,Z),A)\to\mathrm{RHom}(M(X),A)\to\mathrm{RHom}(M(Z),A)$
with $\mathrm{RHom}$ the cochain complex of maps in $\mathit{DM}(\mathrm{Spec}(k))$ (the latter is $\mathbf{Z}$-linear by construction). Since $(a_{!}a^{!}(\mathbf{Z}))^{\wedge}\cong a_{*}a^{*}(\mathbf{Z})$ by Verdier duality, we observe that
$\mathrm{RHom}(M(X),A)\cong\mathrm{RHom}(\mathbf{Z},a_{*}a^{*}(\mathbf{Z}))\cong\mathrm{RHom}(\mathbf{Z}_{X},A_{X})$
where third instance of $\mathrm{RHom}$ is the cochain complex of maps in $\mathit{DM}(X)$, $\mathbf{Z}_{X}=\mathbf{1}_{X}$ is the $\otimes$-unit of $\mathit{DM}(X)$, and $A_{X}=a^{*}(A)$ with $a$ the structural morphism of $X$ as above. Therefore, if $j\colon X-Z\hookrightarrow X$ is the open immersion of the complement of $Z$ in $X$, and if $i\colon Z\hookrightarrow X$ is the embedding of $Z$ into $X$, the cofiber sequence
$j_{!}\,j^{*}(A_{X})\to A_{X}\to i_{*}\,A_{Z}$
thus shows that we have a canonical isomorphism
$\mathrm{RHom}(M(X,Z),A)\cong\mathrm{RHom}(\mathbf{Z}_{X},j_{!}\,j^{*}A_{X})\,.$
Let $h:H\hookrightarrow X$ be another closed embedding. We have a cocartesian square
[…]
hence cartesian squares of the form below.
$\mathrm{RHom}(M(X,Z\cup H),A)$$\mathrm{RHom}(M(X,Z),A)$$\mathrm{RHom}(M(X,H),A)$$\mathrm{RHom}(M(X,H\cap Z),A)$}
Those correspond to cartesian squares
$\mathrm{RHom}(\mathbf{Z}_{X},u_{!}\,u^{*}\,j_{!}\,j^{*}A_{X})$$\mathrm{RHom}(\mathbf{Z}_{X},j_{!}\,j^{*}A_{X})$$\mathrm{RHom}(\mathbf{Z}_{X},u_{!}\,u^{*}A_{X})$$\mathrm{RHom}(\mathbf{Z}_{X},v_{!}\,v^{*}A_{X})$}
where $u\colon X-H\hookrightarrow X$ is the complement of $H$ and $v\colon X-(H\cap Z)\hookrightarrow X$ is the complement of $H\cap Z$ (observe that $u_{!}\,u^{*}A_{X}\cong u_{!}\,u^{*}\,v_{!}\,v^{*}A_{X})$ and $j_{!}\,j^{*}A_{X}\cong j_{!}\,j^{*}\,v_{!}\,v^{*}A_{X})$).
Definition 3.2.
Let $X$ be an affine scheme of finite type over a field $k$ of dimension $d$. A Lefschetz subscheme of $X$ is a closed subscheme $Z$ of $X$ of the form $Z=W\cup H$, where $w\colon W\hookrightarrow X$ is a closed subscheme of dimension <d<d, with affine complement, that contains all irreducible components of $X$ of dimension <d<d, such that XWX-W is regular, and $H$ is an hyperplane section that is transverse to the motivic sheaf $j_{!}\,j^{*}\mathbf{Z}_{X}$ in the sense of 1.18, where $j\colon X-W\to X$ denotes the complement of $W$ in $X$. The rank of the pair $(X,Z)$ is the trace of the identity of M(X,Z)[d]M(X,Z)[-d] in $\mathit{DM}(\mathrm{Spec}(k))$; we will denote it
$\mathit{rk}(X,Z)=(-1)^{d}\mathit{Tr}\big(1_{M(X,Z)}\big)\in\mathbf{Z}$
(this is always an integer because it coincides with the Euler characteristic of its $\ell$-adic realizations [Cis21, Cor. 2.2.7]). A Lefschetz pair is a pair $(X,Z)$ that consists of an affine scheme $X$ and of a Lefschetz subscheme $Z\subset X$.
A Lefchetz filtration of $X$ is a finite sequence of closed subschemes of the form
$\varnothing=X_{-1}\subset X_{0}\subset\ldots\subset X_{d-1}\subset X_{d}=X$
where, for $0\leq i\leq d$, each $X_{i-1}$ is a Lefschetz subscheme of $X_{i}$.
Lefschetz filtrations exist in practice thanks to the following proposition.
Proposition 3.3.
Let $k$ be an infinite field. Then, for any affine $k$-scheme $X$ and any closed subscheme $Y$ in $X$ of dimension $\dim X$, there is a Lefschetz subscheme $Z$ of $X$ containing $Y$. If $k$ is a finite field, this is true only after pulling back $X$ along a finite extension of $k$.
Proof.
Define $W$ to be the union of $Y$ with the singular locus of $X$ as well as with the irreducible components of $X$ of dimension <d<d. We may enlarge $W$ so that its complement if affine, and then apply Corollary 1.17 to $F=j_{!}\,\mathbf{Z}_{X-W}$, where $j\colon X-W\to X$ denotes the complement of $W$ in $X$, in order to find a section $H$ that is transverse to $F$, and put $Z=W\cup H$. The case of a finite field is proved similarly, applying Theorem 1.16 instead of its corollary: the dense open space $U$ of hyperplane sections will only have rational points after a finite extension of $k$. ∎
Lemma 3.4.
Let $X$ be an affine scheme of dimension $d\geq 0$, $Z$ a Lefschetz subscheme of $X$, and $j\colon X-Z\hookrightarrow X$ the corresponding open immersion. Then $\mathit{rk}(X,Z)$ is a non-negative integer. Furthermore, for any Artin-Verdier six-functors formalism $D$ defined on $k$-schemes of finite type, the induced realization functor
$r_{D}\colon\mathit{DM}(\mathrm{Spec}(k))\to D(\mathrm{Spec}(k))$
yields
$H^{q}(X,j_{!}\mathbf{1}_{X-Z})\cong\begin{cases}0&\text{if $q\neq d$,}\\ r_{D}(M(X,Z)[-d])^{\wedge}&\text{if $q=d$.}\end{cases}$
(where the cohomology is defined using the ordinary $t$-structure as in Remark 2.21). The rank of $H^{d}(X,j_{!}\mathbf{1}_{X-Z})$ (defined as the trace of its identity) is equal to the class of $\mathit{rk}(X,Z)$ in $\pi_{0}\operatorname{Map}_{D(\mathrm{Spec}(k))}(\mathbf{1},\mathbf{1})$.
In particular, if $k$ is separably closed, for any prime $\ell$ distinct from the characteristic of $k$, we can compute the $\ell$-adic cohomology of the pair $(X,Z)$ motivically:
$H^{q}_{\mathrm{\acute{e}t}}(X,j_{!}\mathbf{Z}_{\ell})\cong\begin{cases}0&\text{if $q\neq d$,}\\ \mathbf{Z}_{\ell}^{\mathit{rk}(X,Z)}&\text{if $q=d$.}\end{cases}$
Proof.
Let $D$ be an Artin-Verdier six-functors formalism defined on $k$-schemes of finite type. We then have colimit preserving functors
$r_{D}\colon\mathit{DM}(X)\to D(X)$
that commute with the six operations. With the notations of Remark 2.21, if $M(X,Z)^{\wedge}\cong a_{*}j_{!}j^{*}\mathbf{Z}_{X}$ denotes the dual of $M(X,Z)$ in $\mathit{DM}(\mathrm{Spec}(k))$, we thus have
$r_{D}(M(X,Z)^{\wedge})\cong\Gamma(X,j_{!}\,j^{*}\mathbf{1}_{X})\,.$
Therefore,
$r_{D}(M(X,Z)[-d])\cong\Gamma(X,j_{!}\,j^{*}\mathbf{1}_{X}[d])^{\wedge}$
is concentrated in degree 0 by virtue of Theorem 2.22 applied to $M=\mathbf{Z}_{W}$ and $U=W$ (with the notations of Definition 3.2). Since $r_{D}$ is symmetric monoidal, it preserves the formation of traces of endomorphisms of dualizable objects. For any $D$ such that the constructible heart $D^{\heartsuit}_{c}(\mathrm{Spec}(k))$ is tannakian over $\mathbf{Q}$, we get that $\mathit{rk}(X,Z)$ is the Euler characteristic of the image of $a_{*}j_{!}j^{*}\mathbf{Z}_{X}[d]$ through $r_{D}$, and thus a non-negative integer. For instance, if $\ell$ is a prime not equal to the characteristic of $k$, and if $\bar{k}$ is a separable closure of $k$, we can associate to each $k$-scheme $X$ the $\bar{k}$-scheme $\bar{X}$ obtained by pulling back $X$ to $\mathrm{Spec}(\bar{k})$. The assignment
$X\mapsto D(\bar{X},\mathbf{Q}_{\ell})$
then defines an Artin-Verdier six functors formalism $D$ such that $D^{\heartsuit}_{c}(\mathrm{Spec}(k))$ is the category of finite dimensional $\mathbf{Q}_{\ell}$-vector spaces. This proves that $\mathit{rk}(X,Z)$ is a non-negative integer.
Let us assume that $k=\bar{k}$ is separably closed (in order to simplify the notations). We can analyze the situation a little bit further as follows. The morphisms of rings $\mathbf{Q}_{\ell}\leftarrow\mathbf{Z}_{\ell}\to\mathbf{F}_{\ell}$ induce isomorphisms on their Grothendieck rings
$\mathbf{Z}=K_{0}(\mathbf{Q}_{\ell})\cong K_{0}(\mathbf{Z}_{\ell})\cong K_{0}(\mathbf{F}_{\ell})=\mathbf{Z}\,.$
Therefore, for any perfect complex of $\mathbf{Z}_{\ell}$-modules $C$, we always have:
$\sum_{q}(-1)^{q}\dim_{\mathbf{Q}_{\ell}}H^{q}(C)\otimes\mathbf{Q}=\sum_{q}(-1)^{q}\dim_{\mathbf{F}_{\ell}}H^{q}(C\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell})\,.$
If ever both $C\otimes\mathbf{Q}$ and $C\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell}$ are concentrated in cohomological degree $d$, this means that the dimensions of $H^{d}(C\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell})$ and of $H^{d}(C)\otimes\mathbf{Q}$ are the same. A direct consequence can be seen in the exact sequence
$0\to H^{q}(C)\otimes_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell}\to H^{q}(C\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell})\to\mathrm{Tor}_{1}(H^{q+1}(C),\mathbf{F}_{\ell})\to 0\,.$
This yields an isomorphism (an injective linear map from a space that has at least the dimension of the codomain)
$H^{q}(C)\otimes_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell}\cong H^{q}(C\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell})$
for all $q$ (for $q\neq d$ this is even easier) and tells us that $H^{q+1}(C)$ is free for all $q$. Finally, we must have $H^{q}(C)\cong 0$ for any $q\neq d$ and $H^{d}(C)\cong\mathbf{Z}^{r}_{\ell}$ with $r$ the dimension of $H^{d}(C)\otimes\mathbf{Q}$ as a $\mathbf{Q}_{\ell}$-vector space.
On constructible objects, the $\ell$-adic realization functor
$r_{\ell}\colon\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{Z}_{\ell})$
induces the realization functors
$\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{F}_{\ell})\quad\text{and}\quad\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{Z}_{\ell})\otimes\mathbf{Q}=D^{b}_{c}(Y,\mathbf{Q}_{\ell})$
by the formulæ
$M\mapsto r_{\ell}(M)\otimes^{L}_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell}\quad\text{and}\quad M\mapsto r_{\ell}(M)\otimes\mathbf{Q}$
respectively. Therefore, by virtue of Theorem 2.22 for $D^{b}_{c}(-,\mathbf{Q}_{\ell})$ and $D^{b}_{c}(-,\mathbf{F}_{\ell})$, we see that we can apply the discussion above with $C$ be the $\ell$-adic realization of the motive $a_{*}j_{!}j^{*}\mathbf{Z}_{X}[d]$ in the stable $\infty$-category $D^{b}_{c}(\mathrm{Spec}(k),\mathbf{Z}_{\ell})\cong\mathit{Perf}(\mathbf{Z}_{\ell})$. This proves the last part of the lemma. ∎
Remark 3.5.
Let $(X,Z)$ be a Lefschetz pair, with $X$ of dimension $d$ over a separably closed field. Let $j\colon X-Z\to X$ be the corresponding embedding. Given any prime number $\ell$ distinct from the characteristic of the ground field, we have a long exact sequence
$\cdots\to H^{q}_{\mathrm{\acute{e}t}}(X,j_{!}\mathbf{Z}_{\ell})\to H^{q}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})\to H^{q}_{\mathrm{\acute{e}t}}(Z,\mathbf{Z}_{\ell})\to H^{q+1}_{\mathrm{\acute{e}t}}(X,j_{!}\mathbf{Z}_{\ell})\to\cdots$
By virtue of the preceding lemma, this yields isomorphisms
$H^{q}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})\cong H^{q}_{\mathrm{\acute{e}t}}(Z,\mathbf{Z}_{\ell})\quad\text{for $q\neq d,d-1$,}$
and an exact sequence of the form below.
$0\to H^{d-1}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})\to H^{d-1}_{\mathrm{\acute{e}t}}(Z,\mathbf{Z}_{\ell})\to\mathbf{Z}_{\ell}^{\mathit{rk}(X,Z)}\to H^{d}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})\to 0$
The map $H^{d-1}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})\to H^{d-1}_{\mathrm{\acute{e}t}}(Z,\mathbf{Z}_{\ell})$ is thus a split monomorphism inducing an isomorphism on torsion subgroups $H^{d-1}_{\mathrm{\acute{e}t}}(X,\mathbf{Z}_{\ell})_{\mathit{tors}}\cong H^{d-1}_{\mathrm{\acute{e}t}}(Z,\mathbf{Z}_{\ell})_{\mathit{tors}}$.
Theorem 3.6.
Let $X$ be an affine scheme of Krull dimension $d$, of finite type over a separably closed field $k$, and equipped with a Lefschetz filtration
$\varnothing=X_{-1}\subset X_{0}\subset\ldots\subset X_{d-1}\subset X_{d}=X\,.$
For any prime $\ell$ distinct from the characteristic of $k$, and $0\leq q\leq d$, we have:
$\mathit{rk}\,H^{q}(X,\mathbf{Z}_{\ell})\leq\mathit{rk}\,H^{q}(X_{q},\mathbf{Z}_{\ell})\leq\mathit{rk}(X_{q},X_{q-1})\,.$
Proof.
This follows right away from Lemma 3.4 and from Remark 3.5 by induction on $d$. ∎
Remark 3.7.
Together with Proposition 3.3, this gives uniform bounds of Betti numbers for all quasi-projective schemes over a field, using Jouanolou’s trick [Jou73, Lemme 1.5]: the latter proves that any quasi-projective scheme is $\mathbf{A}^{1}$-homotopy equivalent to an affine one.
Lefschetz pairs may be used to prove the existence of uniform bounds of Betti numbers of any constructible motive; see Corollary 3.10 below.
3.8.
We can interpret the existence of Lefschetz filtrations as an approximation of the motivic $t$-structure on $\mathit{DM}(X)$ (whose existence would imply Grothendieck’s standard conjectures [Bei10]). In order to simplify our task, we will work with $\mathbf{Q}$-linear coefficients (it is straightforward to produce an integral version from any $\mathbf{Q}$-linear one using the rigidity theorem).
We observe that objects of the form $M(X)(n)$, with $X$ affine of finite type over the base field $k$ and $n\in\mathbf{Z}$ are compact and generate $\mathit{DM}(\mathrm{Spec}(k))$ as a compactly generated stable $\infty$-category. If $k$ is infinite, this implies that the objects of the form $M(X,Z)(n)$ with $(X,Z)$ a Lefschetz pair and $n\in\mathbf{Z}$ also form a generating family. In order to allow any field, we proceed as follows.
Proposition 3.9.
Let $k$ be any field. We define the family $\mathcal{G}_{k}$ as follows: this is the family of motives of the form $f_{*}M(X,Z)[-\dim X](n)$, where $f\colon\mathrm{Spec}(K)\to\mathrm{Spec}(k)$ is a Galois extension of the base field, with $(X,Z)$ a Lefschetz pair defined over $K$, and $n\in\mathbf{Z}$. Then $\mathcal{G}_{k}$ is a family of compact generators of $\mathit{DM}(k,\mathbf{Q})$ as a cocomplete stable $\infty$-category.
Proof.
The case where $k$ is infinite is obvious from Proposition 3.3: any affine scheme has a Lefschetz filtration and given a Lefschetz filtration by subschemes $X_{i}$, one can reconstruct the motive of $X$ from the $M(X_{i},X_{i-1})$’s. Since the motives of the form $M(X)(n)$ with $X$ affine form a generating family, this proves the claim if $k$ is infinite, in particular for $k$ separably closed. In general, we know that, if $\bar{k}$ is a separable closure of $k$, then the pullback functor
$\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})\to\mathit{DM}(\mathrm{Spec}(\bar{k}),\mathbf{Q})$
is conservative and preserve the formation of internal Hom’s of the form $\mathit{Hom}(A,B)$ as long as $A$ is compact; see [CD19, Lemma 4.3.7]. By a continuity argument (as in [CD19, Prop. 4.3.1]), it is now easy to see that it suffices to prove that $\mathcal{G}_{k}$ generates $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$ if $\mathcal{G}_{\bar{k}}$ generates $\mathit{DM}(\mathrm{Spec}(\bar{k}),\mathbf{Q})$, which we already know. ∎
Corollary 3.10.
Let $M$ be any constructible object in $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$. There exists a constant $C>0$ and integer $a\leq b$ with the following properties. For any $\mathbf{Q}$-linear Artin-Verdier six-functors formalism $D$ defined over $k$-schemes of finite type with tannakian heart $D^{\heartsuit}_{c}(\mathrm{Spec}(k))$, the induced functor
$r_{D}\colon\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})\to D(\mathrm{Spec}(k))$
sends $M$ to an object $r_{D}(M)$ of cohomological amplitude $[a,b]$ with respect to the ordinary $t$-structure and so that the rank of each cohomology objects $H^{q}(r_{D}(M))$ (i.e. the traces of their identity) is bounded by $C$.
Proof.
The preceding proposition means that such an $M$ is obtained by finitely many operations (such as extracting a direct factor in a finite coproducts, taking cofibers, shifting by some integer) from finitely many objects of the form $f_{*}M(X,Z)[-\dim X](n)$, where $f\colon\mathrm{Spec}(K)\to\mathrm{Spec}(k)$ is a Galois extension of the base field, with $(X,Z)$ a Lefschetz pair defined over $K$, and $n\in\mathbf{Z}$. Therefore, it suffices to prove the result for $M=f_{*}M(X,Z)[-\dim X](n)$. Since $f_{*}$ is an exact functor for $f$ finite, we conclude with Lemma 3.4. ∎
3.11.
We call standard generators objects of the form
$A^{\wedge}_{1}\otimes\cdots\otimes A^{\wedge}_{m}\otimes B_{1}\otimes\cdots\otimes B_{n}$
where each $A_{i}$ and $B_{j}$ are in $\mathcal{G}_{k}$ and $(-)^{\wedge}$ is the duality operator on dualizable objects in $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$. We define $\mathit{DM}^{\leq 0}(\mathrm{Spec}(k),\mathbf{Q})$ as the smallest full subcategory of $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$ stable under small colimits and stable under extensions spanned by standard generators. This is a compactly generated presentable $\infty$-category. Therefore, the inclusion functor
$\mathit{DM}^{\leq 0}(\mathrm{Spec}(k),\mathbf{Q})\hookrightarrow\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$
has a right adjoint
$\tau^{\leq 0}\colon\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})\to\mathit{DM}^{\leq 0}(\mathrm{Spec}(k),\mathbf{Q})$
We define $\mathit{DM}^{\geq 0}(\mathrm{Spec}(k),\mathbf{Q})$ as the full subcategory of objects $B$ such that
$\operatorname{Map}(A[1],B)=0$
for any $A$ in $\mathit{DM}^{\leq 0}(\mathrm{Spec}(k),\mathbf{Q})$. This defines a $t$-structure on $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$. By construction (and thanks to Lemma 3.4), for any $\mathbf{Q}$-linear Artin-Verdier six-functors formalism $D$ defined over $k$-schemes of finite type, the induced functor
$r_{D}\colon\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})\to D(\mathrm{Spec}(k))$
is right $t$-exact and induces a symmetric monoidal colimit preserving functor
$r_{D}\colon\mathit{DM}^{\leq 0}(\mathrm{Spec}(k),\mathbf{Q})\to D^{\leq 0}(\mathrm{Spec}(k))\,.$
This $t$-structure on $\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$ is an approximation of the conjectural motivic $t$-structure - if we are optimistic, we could hope it is the right one. We would like to do something similar in $\mathit{DM}(Y,\mathbf{Q})$ for any $Y$ of finite type over a field: approximating the perverse $t$-structure. For this, we need to define a relative notion of Lefschetz pairs.
3.12.
Let $Y$ be a regular scheme of finite type over $k$. An affine $Y$-scheme of finite type $X$ is admissible if the structural map $a\colon X\to Y$ is flat and if the motivic sheaf $a_{*}(\mathbf{Q}_{X})$ is dualizable in $\mathit{DM}(Y,\mathbf{Q})$ (up to replacing $Y$ by a dense open subscheme, any flat scheme over $Y$ is admissible; by Grothendieck’s generic flatness theorem, this means that, up to replacing $Y$ by a dense open subscheme, any $Y$-scheme of finite type can be made admissible).
Given an admissible $Y$-scheme $X$ with structural morphism $a\colon X\to Y$, we define
$M_{Y}(X)=a_{*}(\mathbf{Z}_{X})^{\wedge}\,.$
Given a closed subscheme $Z$ of $X$ that is admissble over $Y$ as well, we define $M_{Y}(X,Z)$ by a cofiber sequence
$M_{Y}(Z)\to M_{Y}(X)\to M_{Y}(X,Z)$
Definition 3.13.
Let $Y$ be a regular scheme of finite type over $k$. Let $a\colon X\to Y$ be flat morphism between $k$-schemes of finite type of relative dimension $d$. A Lefschetz subscheme of $X$ over $Y$ is a closed subscheme $Z$ of $X$ that is flat over $Y$, of the form $W\cup H$, where $W$ is a closed subscheme of $X$, with affine complement, and that if flat over $Y$, with $\dim(W)=\dim(X)-1$, which contains all irreducible components of $X$ of dimension $\dim(X)$, such that XWX-W is regular, and $H$ is an hyperplane section that is transverse to the motivic sheaf $j_{!}\,j^{*}\mathbf{Z}_{X}$ in the sense of 1.18, where $j\colon X-W\to X$ is the embedding of the complement of $W$. We will then say that $(X,Z)$ is a Lefschetz pair over $Y$. A Lefschetz pair $(X,Z)$ is admissible over $Y$ if both $X$ and $Z$ are admissible.
3.14.
Let $a\colon X\to Y$ be the structural morphism of the admissible affine $Y$-scheme $X$. We choose an affine open embedding $e\colon X\to\bar{X}$ followed by a projective morphism $\bar{a}\colon\bar{X}\to Y$ with $a=\bar{a}\,e$. Possibly after shrinking $Y$, we may assume that $\bar{X}$ is flat over $Y$. Let us consider an admissible Lefschetz subscheme $Z$ in $X$, and write $p\colon\bar{X}-\bar{H}\to Y$ for the structural map of the complement of the chosen hyperplane section $H$. We have
$a_{*}\,j_{!}\,\mathbf{Q}_{X-Z}\cong p_{!}\,e_{*}\,j_{!}\mathbf{Q}_{X-W}\,.$
We will say that $(X,Z)$ is tight if the conclusion of Proposition 2.20 holds without shrinking $Y$ any further for the realizations of $a_{*}\,j_{!}\,\mathbf{Q}_{X-Z}$ and of $p_{!}\,e_{*}\,j_{!}\mathbf{Q}_{X-W}$ through any $\mathbf{Q}$-linear Artin-Verdier six-functors formalism $D$. The aforementioned proposition says that, given any admissible Lefschetz pair $(X,Z)$ over $Y$, there always exists a dense open subscheme of $Y$ over which $(X,Z)$ becomes tight.
Proposition 3.15.
If $(X,Z)$ is a tight admissible Lefschetz pair over a regular scheme $Y$, with $X$ of relative dimension $d$ over $Y$, then, for any $\mathbf{Q}$-linear Artin-Verdier six-functors formalism $D$, the induced functor
$r_{D}\colon\mathit{DM}(Y,\mathbf{Q})\to D(Y)$
sends
$M_{Y}(X,Z)^{\wedge}[d]\cong a_{*}\,j_{!}\,\mathbf{Q}_{X-Z}[d]$
to a dualizable object that is concentrated in degree 0 for the ordinary $t$-structure in $D(Y)$.
Proof.
This is a direct consequence of the definition, exactly as in the proof of Theorem 2.22 and of Lemma 3.4. ∎
3.16.
Let $Y$ be a scheme of finite type over $k$ with $Y_{\mathit{red}}$ regular, and let $\imath\colon Y_{\mathit{red}}\to Y$ be the canonical embedding. We denote by $\mathcal{G}_{Y}$ the family of objects of the form
$\imath_{*}\,f_{*}M_{Y^{\prime}}(X^{\prime},Z^{\prime})(n)[-d]$
with $f\colon Y^{\prime}\to Y_{\mathit{red}}$ a finite flat morphism, $Y^{\prime}$ regular, $(X^{\prime},Z^{\prime})$ a tight admissible Lefschetz pair over $Y^{\prime}$, $d$ the relative dimension of $X$ over $Y^{\prime}$ (seen as a locally constant function), and $n$ any integer. As we did previously we define standard generators over $Y$ objects of the form
$A^{\wedge}_{1}\otimes\cdots\otimes A^{\wedge}_{m}\otimes B_{1}\otimes\cdots\otimes B_{n}$
where each $A_{i}$ and $B_{j}$ are in $\mathcal{G}_{Y}$ and $(-)^{\wedge}$ is the duality operator on dualizable objects in $\mathit{DM}(Y,\mathbf{Q})$.
For a scheme of finite type $X$ over $k$, we define the perverse generating family $P(X)$ by noetherian induction on dimension as follows. For $\dim X\leq 0$, $P(X)=\mathcal{G}_{X}$. For $\dim X>0$, $P(X)$ is the collections of objects of the following two possible forms:
  • $i_{*}(M)$ for $i\colon Z\to X$ a nowhere dense closed immersion and $M$ in $P(Z)$;
  • $j_{!}M[\dim Y]$ for $j\colon Y\to X$ an affine dense open immersion with $Y_{\mathit{red}}$ affine and regular, and $M$ is a standard generator over $Y$.
Proposition 3.17.
The family $P(X)$ forms a family of compact generators in $\mathit{DM}(X,\mathbf{Q})$.
Proof.
We already know this is true in the case where $X$ is the spectrum of a field. The case where $X$ is of dimension $\leq 0$ follows right away. Theorem 1.16 together with Proposition 2.20 and Grothendieck’s generic flatness theorem ensure that tight admissible Lefschetz pairs exist in abundance generically. The proposition follows easily from [CD19, Prop. 4.3.17] applied to $\mathit{DM}$ by continuity arguments. ∎
3.18.
We define $^{p}\mathit{DM}^{\leq 0}(X,\mathbf{Q})$ as the smallest full subcategory of $\mathit{DM}(X,\mathbf{Q})$ stable under small colimits and stable under extensions spanned by $P(X)$. This is a compactly generated presentable $\infty$-category. Therefore, the inclusion functor
$^{p}\mathit{DM}^{\leq 0}(X,\mathbf{Q})\hookrightarrow\mathit{DM}(\mathrm{Spec}(k),\mathbf{Q})$
has a right adjoint
$\tau^{\leq 0}\colon\mathit{DM}(X,\mathbf{Q})\to{{}^{p}\mathit{DM}^{\leq 0}(X,\mathbf{Q})}$
We define $^{p}\mathit{DM}^{\geq 0}(X,\mathbf{Q})$ by orthogonality as the full subcategory of objects $B$ such that
$\operatorname{Map}(A[1],B)=0$
for any $A$ in $^{p}\mathit{DM}^{\leq 0}(X,\mathbf{Q})$. This defines a $t$-structure on $\mathit{DM}(X,\mathbf{Q})$ that would be a candidate for the perverse motivic $t$-structure.
If $D$ is any $\mathbf{Q}$-linear Artin-Verdier six-functors formalism defined over $k$-schemes of finite type, the perverse $t$-structure on $D_{c}$ extends to a perverse $t$-structure on $D$ such that the embedding
$D_{c}(X)\hookrightarrow D(X)$
is $t$-exact for all $X$: simply define $^{p}D^{\leq 0}(X)$ as the smallest full subcategory of $D(X)$ stable under small colimits and stable under extensions spanned by perverse sheaves in $D_{c}(X)$. For any $k$-scheme of finite type $X$, Proposition 3.15 imply that the induced functor
$r_{D}\colon\mathit{DM}(X,\mathbf{Q})\to D(X)$
is right $t$-exact and induces a colimit preserving functor
$r_{D}\colon{{}^{p}\mathit{DM}^{\leq 0}(X,\mathbf{Q})}\to{{}^{p}D^{\leq 0}(X)}\,.$
There is an analogous construction for a candidate of the ordinary $t$-structure on $\mathit{DM}(X,\mathbf{Q})$. Define $G(X)$ as follows: $G(X)=P(X)$ for $X$ of dimension $\leq 0$, and for $X$ of dimension $d>0$, define $G(X)$ as the collection of objecrs of the form
  • $i_{*}(M)$ for $i\colon Z\to X$ a nowhere dense closed immersion and $M$ in $G(Z)$;
  • $j_{!}M$ for $j\colon Y\to X$ an affine dense open immersion with $Y_{\mathit{red}}$ affine and regular, and $M$ is a standard generator over $Y$.
We define $\mathit{DM}^{\leq 0}(X,\mathbf{Q})$ as the smallest full subcategory of $\mathit{DM}(X,\mathbf{Q})$ stable under small colimits and stable under extensions spanned by $P(X)$. This defines as above a $t$-structure such that $r_{D}$ is right $t$-exact for any $D$ (with respect to the ordinary $t$-structure on $D$). By construction, $\mathbf{Q}_{X}$ belongs to $\mathit{DM}^{\leq 0}(X,\mathbf{Q})$ and the tensor product defines a functor that preserves colimit in each variables
$(-)\otimes(-)\colon\mathit{DM}^{\leq 0}(X,\mathbf{Q})^{\mathit{op}}\times\mathit{DM}^{\leq 0}(X,\mathbf{Q})\to\mathit{DM}^{\leq 0}(X,\mathbf{Q})\,.$
A wild optimistic guess would be that any element of $G(X)$ is in the heart of this $t$-structure.